Nicole:

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5).

Nicole:

for the previous question, that was point slope form right?

AngeI:

Slope is -1 so Y=-1x+b Add in (-2,5) for x and y 5=-1(-2)+b Getting b=3 Y=-1x+3

Nicole:

but it says we need point slope form not slope intercept

AngeI:

You can convert it o: I assumed you knew how How much time do you have left so I can see if I can explain it?

Nicole:

im not positive can you help me please?

AngeI:

Y=-1x+3 You would do (-2,5) for x1 and y1 Y-5=-1(x-(-2)) Took me a bit I had to walk to my different class